OpenStudy (anonymous):

WILL MEDAL: -Linear Equation- Write the following in standard from: y-3 = (-1/4)(x+4). I got: x + 4y = 4. Please double check my answer. Thanks.

OpenStudy (calculusfunctions):

The right side of your final answer is incorrect. It's not 4.

OpenStudy (anonymous):

y-3 = (-1/4)(x+4) y = -(1/4)x + (4/4) + 3 Since (-1/4) = -(1/4); then -(1/4)(4) = +(4/4)?

OpenStudy (anonymous):

y = -(1/4)x + (4/4) + 3 y = -(1/4)x + (16/4) x + 4y = 16?

OpenStudy (anonymous):

y = -(1/4)x + (-4/4) + 3

OpenStudy (calculusfunctions):

No, in your former response, when you expanded, you should have y = -(1/4)x - (4/4) + 3

OpenStudy (anonymous):

I'm getting my signs fudge up again So: x + 4y = -1?

OpenStudy (anonymous):

Damn it -8?

OpenStudy (anonymous):

y = -(1/4)x + (-1) + 3 y = -(1/4)x + 2

OpenStudy (calculusfunctions):

No, look\[y -3=\frac{ -1 }{ 4 }(x +4)\]\[y =-\frac{ 1 }{ 4 }x -1+3\]\[y =-\frac{ 1 }{ 4 }x +2\]Now multiplying both sides by 4 yields\[4y =-x +8\]Thus\[x +4y =8\]

OpenStudy (calculusfunctions):

So you see your error.

OpenStudy (calculusfunctions):

Anything else I can assist you with?

OpenStudy (anonymous):

Yes thanks. I'll just post another question if I need help in anything else.